OpenStudy (chris215):

Find all solutions in the interval [0, 2π). 7 tan3x - 21 tan x = 0

OpenStudy (chris215):

is this the answer?

OpenStudy (darthvader2900):

http://openstudy.com/study#/updates/50f1ca0ce4b0694eaccf36a0 there ya go

OpenStudy (freckles):

so you have \[7 \cdot \tan(3x)-21 \cdot \tan(x)=0\] We will probably have to attempt to write tan(3x) in terms of tan(x) somehow using identities

OpenStudy (freckles):

like if possible that is

OpenStudy (freckles):

we can start out first with dividing both sides by 7 \[\tan(3x)-3 \tan(x)=0\]

OpenStudy (freckles):

My first thought it to rewrite tan(3x)=tan(2x+x) using the sum identity for tan

OpenStudy (chris215):

how did you get to tan(3x)=tan(2x+x)

OpenStudy (freckles):

2x+x is 3x

OpenStudy (chris215):

oh ok

OpenStudy (freckles):

use the sum formula for tan(2x+x)

OpenStudy (freckles):

then use the double angle formula for tan(2x)

OpenStudy (freckles):

before we continue are you sure you didn't mean to write 7*tan^3(x)-21tan(x)=0?

OpenStudy (chris215):

yeah that's the right equation sorry

OpenStudy (freckles):

oh that makes thing a lot more easier

OpenStudy (freckles):

\[7 \tan^3(x)-21\tan(x)=0\\ \text{ divide both sides by } 7 \\ \tan ^3(x)-3 \tan(x)=0\] now just factor the left hand side

OpenStudy (freckles):

notice that both terms have a tan(x) in common

OpenStudy (freckles):

have you done that yet ? if so can you type here please ?

OpenStudy (chris215):

tanx = 0 or tan^2 x = 3 ?

OpenStudy (freckles):

let's see you have \[\tan(x)(\tan^2(x)-3)=0 \\ \tan(x)=0 \text{ or } \tan^2(x)=3 \\ \tan(x)=0 \text{ or } \tan(x)=\sqrt{3} \text{ or } \tan(x)=-\sqrt{3}\] you have 3 equations to solve in the interval [0,2pi]

OpenStudy (freckles):

tan(x)=sqrt(3) happens when sin(x)=sqrt(3)/2 and cos(x)=1/2 or when sin(x)=-sqrt(3)/2 and cos(x)=-1/2 (you can find when tan(x)=-sqrt(3) in a similar way; just make sin and cos differ in sign) tan(x)=0 happens when sin(x)=0

OpenStudy (freckles):

can you solve sin(x)=0?

OpenStudy (chris215):

how do I do that

OpenStudy (freckles):

use the unit circle

OpenStudy (freckles):

https://www.mathsisfun.com/geometry/images/circle-unit-radians.gif look for which angles the y-coordinate is 0

OpenStudy (freckles):

for example you see the angle 0 has y-coordinate 0 there is another angle so that the y-coordinate is 0, what is it?

OpenStudy (chris215):

pi/2?

OpenStudy (freckles):

at pi/2 the y-coordinate is 1

OpenStudy (chris215):

0?

OpenStudy (freckles):

the y-coordinate is the second number in the ordered pairs

OpenStudy (freckles):

I already said at the angle 0 the y-coordinate is 0 I'm asking you to give me another answer looking at the unit circle

OpenStudy (chris215):

another angle that the y-coordinate is 0? I see two points (1,0) and (-1,0) anf pi and 0

OpenStudy (freckles):

yes! bingo!

OpenStudy (freckles):

tan(x)=0 has solutions x=0 or x=pi

OpenStudy (freckles):

now you have to solve tan(x)=sqrt(3) then solve solve tan(x)=-sqrt(3)

OpenStudy (freckles):

tan(x)=sqrt(3) only happens when tan is positive which only happens when sin and cos have the same signs so look in the first and third quadrant look for when the sin value there is either sqrt(3)/2 or -sqrt(3)/2

OpenStudy (freckles):

for tan(x)=-sqrt(3) this happens only in the 2nd and fourth quadrants (you know where sin and cos have opposite signs) so here you would look for where again sin value is either sqrt(3)/2 or -sqrt(3)/2

OpenStudy (freckles):

the sin value being the y-coordinates

OpenStudy (chris215):

ok so are all those the solutions

OpenStudy (freckles):

well you only found two solutions there are 4 more

OpenStudy (chris215):

none of those are choices though

OpenStudy (freckles):

I see 0,pi you haven't even attempted to find the other four solutions

OpenStudy (freckles):

I told you how above but I'm not sure why you haven't tried to

OpenStudy (freckles):

for example tan(x)=sqrt(3) the hint I gave to find your solutions to this was to look in the first and third quadrant and find when the sin value was either sqrt(3)/2 or -sqrt(3)/2 that happens at the angles pi/3 and 4pi/3

OpenStudy (freckles):

you try to solve tan(x)=-sqrt(3)

OpenStudy (chris215):

tanx = tan (π/6) or tan (π + π/6)

OpenStudy (freckles):

Is this the next question?

OpenStudy (freckles):

Or are you trying to solve tan(x)=-sqrt(3)?

OpenStudy (freckles):

Ok the sin value is the y-coordinate do you see anywhere in the 2nd or 4th quadrant that the y-coordinate is sqrt(3)/2 or -sqrt(3)/2

OpenStudy (chris215):

yes

OpenStudy (freckles):

at what angles?

OpenStudy (freckles):

Do you see the angle for which the y-coordinate is sqrt(3)/2 in the 2nd quadrant? Do you see the angle for which the y-coordinate is -sqrt(3)/2 in the 4th quadrant?

OpenStudy (freckles):

\[\text{ on } [0,2 \pi] \\ \tan(x)=0 \text{ when } x=0,\pi \\ \tan(x)=\sqrt{3} \text{ when } x=\frac{\pi}{3}, \frac{4 \pi}{3} \\ \tan(x)=-\sqrt{3} \text{ when } x=\frac{2 \pi}{3}, \frac{5\pi}{3}\]

OpenStudy (chris215):

ok then what

OpenStudy (freckles):

I don't think we can do anything further :p

OpenStudy (freckles):

What do you want to do further? Besides solve the equation?

OpenStudy (chris215):

oh ok thank you so much!

OpenStudy (freckles):

Like all I seen where choices to the equation above which I already found for you

OpenStudy (freckles):

I think you really need to work on learning how to read the unit circle.

OpenStudy (freckles):

like for example sin(x)=1/2 means look for the angles where you have the y-coordinate is 1/2

OpenStudy (freckles):

so you go to your unit circle

OpenStudy (chris215):

yeah I think I have somewhat of an understanding now

OpenStudy (freckles):

and you see it happens at pi/6 or also 5pi/6

OpenStudy (chris215):

yeah

OpenStudy (freckles):

so the solutions to sin(x)=1/2 is therefore x=pi/6 or 5pi/6

OpenStudy (freckles):

that had nothing to do with your question it was just another example

OpenStudy (chris215):

thanks that was really helpful

OpenStudy (freckles):

np have a lovely day